OpenStudy (anonymous):

Solve, using the substitution method. y = 2x + 6 3x – y = 6 A. There are an infinite number of solutions. B. (12, 30) C. (14, 12) D. There is no solution.

OpenStudy (anonymous):

@PigPresident

OpenStudy (anonymous):

12

OpenStudy (anonymous):

is it b or c

OpenStudy (anonymous):

y = 2x + 6 now sub 2x + 6 in for y in the other equation 3x - y = 6 3x - (2x + 6) = 6 3x - 2x - 6 = 6 3x - 2x = 6 + 6 x = 12 now sub 12 in for y in either equation to find y y = 2x + 6 y = 2(12) + 6 y = 24 + 6 y = 30 check... 3x - y = 6 3(12) - 30 = 6 36 - 30 = 6 6 = 6 (correct) solution is : (12,30)

OpenStudy (anonymous):

o ok thanks

OpenStudy (anonymous):

typo...sub 12 in for x your welcome..glad to help :)